OpenStudy (yb1996):

Find the smallest possible value of the constant A such that ln(x) <= Ax^2 for all x>0.

ganeshie8 (ganeshie8):

Hey

OpenStudy (yb1996):

Hello

ganeshie8 (ganeshie8):

tried anything ?

OpenStudy (yb1996):

No, I'm not really sure what I'm supposed to do to be able to get the answer

ganeshie8 (ganeshie8):

Have you understood the question ?

ganeshie8 (ganeshie8):

Notice that `lnx` grows slowly compared to `x`

ganeshie8 (ganeshie8):

Honestly i also don't know how to solve this problem, im still thinking..

OpenStudy (yb1996):

Yes, so I'm guessing the question is asking for A, so that x^2 can just touch the graph of ln(x)

OpenStudy (sparrow2):

i think derivatis must be equal : so 1/x=2ax a=1/2x^2

OpenStudy (whpalmer4):

Well, numerically, the value of A is bounded by \[0.183935 < A < 0.18394\]

OpenStudy (whpalmer4):

narrowed it further: \[0.18393972058572 < A < 0.18393972058573\] Should be a simple matter to determine the symbolic version, right? :-)

Parth (parthkohli):

One thing is definitely true: when they meet, this particular curve will rise above faster than the other one.Let's say they meet at some value \(x=t\) where \(t > 0\). Their tangents must coincide there, and so their slopes must coincide.\[\Rightarrow \frac{1}t = 2At\]\[\Rightarrow 2At^2 = 1 \Rightarrow A = \frac{1}{2t^2}\]Also \(\ln t = At^2\) therefore \(\ln t = \frac{1}2\Rightarrow t = e^{1/2}\). Now our curve \(y = Ax^2\) passes through \((e^{1/2},1/2)\) and so \(A = \frac{1}{2e} \).